Assume that you contribute $100 per month to a retirement plan for 20 years. Then you are able to increase the contribution to $200 per month for another 20 years. Given a 6 percent interest rate, what is the value of your retirement plan after 40 years?
Answer #1

N = 40 × 12 = 480, I = 6/12 = 0.5, PV = 0, PMT = 100, CPT FV = 199149
N = 20 × 12 = 240, I = 6/12 = 0.5, PV = 0, PMT = 100 (200 − 100), CPT FV = 46204
sum of FV = 245353.

answered by: Kraer4342